OpenStudy (anonymous):

Solve for x: 2/3 (x - 2) = 4x. A. -2/5 B. -5/2 C. 2/5 D. 5

Elsa213 (elsa213):

\(x^2+4x=0 x(x+4)=0 so, x=0 and x=-4 w.\)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

? I don't understand what u just put

OpenStudy (anonymous):

\[x - 2= 4x\\ -2=3x\\ -\frac{2}{3}=x\]

OpenStudy (anonymous):

that is because it was wrong subtract \(x\) from both sides, then divide by \(3\)

OpenStudy (anonymous):

X would be negative 2/3

OpenStudy (anonymous):

but thats not one of the options.

Elsa213 (elsa213):

what r the options?

OpenStudy (anonymous):

I think its A but not for sure what lesson is this and what school.

OpenStudy (anonymous):

A. -2/5 B. -5/2 C. 2/5 D. 5

OpenStudy (anonymous):

and am taking algebra on online school

OpenStudy (anonymous):

Be sure that both the question and the options for answers are written correctly.

OpenStudy (anonymous):

the question was written wrong didn't put the 2/3 but the answers are correct.

OpenStudy (anonymous):

So can you rewrite the question?

OpenStudy (anonymous):

Solve for x: 2/3 (x - 2) = 4x

OpenStudy (anonymous):

I think its A so am just going to put that

OpenStudy (anonymous):

give me a medal for it